OpenStudy (anonymous):

The curve is given as y = ln(x) - 1. Find the equation of the tangent to this curve that is parallel to the line 3x-6y-1 = 0. Ok, so i differientiated y = ln(x) - 1 and found x = -1/2 Something is not right here... plz helpp

OpenStudy (anonymous):

\[\frac{1}{x}=\frac{1}{2}\] solve for \(x\) in your head

OpenStudy (anonymous):

x = 2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

slope is \(\frac{1}{2}\), point is \((2,\ln(2)-1)\)

OpenStudy (anonymous):

what? but i got x = -1/2... the derivative is dy/dx = 1/x and i set dy/dx = -2 and solved for x

OpenStudy (anonymous):

cause i thought the slope was parallel to the other line

OpenStudy (anonymous):

\[3x-6y-1 = 0\] what is the slopoe?

OpenStudy (anonymous):

oh, nvm, i see. I thought it was perpendicular

OpenStudy (anonymous):

k

OpenStudy (anonymous):

thank you !

OpenStudy (anonymous):

yw